Question 14 O pts Find the equation of the parabola with focus (10, -3) and directrix y = 3.
Answer #1

9 -3 (0) + (95) 2 2. د (92) = 2 ( 3+ و + (10-1) ++ و - و = 7 + 9 ) + کو 2 + 5) + لاہور - 12 ㅋ = و 2 + SD + 20 - 12 - (1-10) +